OpenStudy (anonymous):

3(2x+3(x-4x))-5x+1 Is the answer -44x+1? if not where did i go wrong? I Started with x-4x and got -5x then i did 3 multiplied by -5x and got -15x. then i did 2x+-15x and got -13x. then i did 3 multiplied by -13x and got -39x. then i did -39x-5x and got -44x+1

OpenStudy (anonymous):

To start off, x-4x is -3x, not -5x

OpenStudy (anonymous):

so that is where you first went wrong

OpenStudy (anonymous):

3(2x+3x-12x)-5x+1 6x+9x-36x-5x+1 -26x+1

OpenStudy (anonymous):

because: \[1x-4x = (1-4)x = -3x\] using the distributive property

OpenStudy (anonymous):

@greysica gives you the correct result

OpenStudy (anonymous):

Thanks sooo much!

OpenStudy (anonymous):

you are welcome